OpenStudy (anonymous):

Find the slopes of lineRSand lineTV. Then determine whether lineRSand lineTV are parallel, perpendicular, or neither. 7. R(-2, 6), S(3, 4), T(3, 5), V(0,0) (1 point) Slope of lineRS = -twofifths_large; slope of lineTV = fivethirds_large; parallel Slope of lineRS = -twofifths_large; slope of lineTV = fivethirds_large; perpendicular Slope of lineRS = -twofifths_large; slope of lineTV = threefifths_large; neither Slope of lineRS = -twofifths_large; slope of lineTV = fivethirds_large; neither

OpenStudy (jit4won):

To find the slope you should use slope(m)= change in y coordinate/change in x coordinate

OpenStudy (jit4won):

slope of RS= ( 4-6)/(3+2)=-2/5 slope of TV= (0-5)/(0-3)=5/3

OpenStudy (jit4won):

Slope of lineRS = -twofifths_large; slope of lineTV = fivethirds_large; neither

OpenStudy (jit4won):

they can't be parallel as their slopes are not equal and; they can't be perpendicular as product of their slope is not equal to -1

OpenStudy (anonymous):

so whats the answer? a b or c? thx

OpenStudy (anonymous):

nevermind thanyou :)